find the equation of an ellipse


Centre is at orgin,major axis is along x axis,e=2/3 and passess throught the point (2,-5/3)

Grade:12

1 Answers

Vikas TU
11 years ago

1. let eqn. of the ellipse be x2/a2 + y2/b2 = 1...........................($)

Now pass the points    as it satifies the ellipse!

You will get a eqn. in the form of a and b

2. Now, as we know that

b2=a2(e2-1).............................(@)

eccentricity is given = 2/3

put in the eqn. (@) u will get an another eqn. in the form of a and b

3. nOw u have two variables and two eqns. solve them and get the values kif a  and b!

4. Put the values of a and b into the eqn. of the ellipse

eqn............($)

and get the eqn of the ellipse!

 

Hope it helped u!